OpenStudy (anonymous):

Please help, will give medal and fan. Algebra 1

OpenStudy (anonymous):

What is the y-intercept of the function, and what does this tell you about the turtle? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 4 hours, and tell what the average rate represents.

OpenStudy (animelover98):

The y tells you how far away the turtle is from the nest in miles.

OpenStudy (anonymous):

The y-intercept is (0,5)

OpenStudy (anonymous):

and it says that the turtle started 5 miles from its nest?

OpenStudy (animelover98):

Yes

OpenStudy (animelover98):

So at point (0,5) the turtle started.

OpenStudy (anonymous):

when I calculated the rate of change I got \[\frac{ 66 }{ 3 }\] I don't know if I should simplify that or not...

OpenStudy (animelover98):

When calculating you used rise/run correct?

OpenStudy (anonymous):

no I used\[\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (anonymous):

which is basically the same thing, wow, so yes

OpenStudy (animelover98):

Ok good

OpenStudy (anonymous):

so should I simplify?

OpenStudy (animelover98):

So, you would want to simplify

OpenStudy (animelover98):

Just like if you had gotten 8/4, you would simplify it to 2.

OpenStudy (anonymous):

and it represents how fast the turtle is going?

OpenStudy (animelover98):

Yes, it represents the rate of change between the two measurement.

OpenStudy (animelover98):

Which would be the speed of the turtle

OpenStudy (anonymous):

ok. can you help me with one more thing?

OpenStudy (animelover98):

Sure

OpenStudy (anonymous):

What would be the domain of the function if the turtle continued to walk at this rate until it traveled 225 miles from the nest? To get to 225 the x value is 15. For the domain should I write out 1,2,3,4,5,6,7,8,9,10,11,12,13,14,15 or just make an inequality or something?

OpenStudy (animelover98):

You would make the inequality

OpenStudy (anonymous):

so it would look like\[1\le x \le15\]

OpenStudy (animelover98):

No I'm sorry! Not the inequality

OpenStudy (animelover98):

http://www.purplemath.com/modules/fcns2.htm

OpenStudy (animelover98):

I have to log off, but this article should help you. There is also a video.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (animelover98):

Of course! I hope I helped